# What Is That Talk About? A Video-to-Text Summarization Dataset for Scientific Presentations

### v3

@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 948878).
@@ Appendix D Model Version Details
LLaMA-VID Li et al. (2025) LLaMA-VID-7B-Full-224-Long-Video 7B
LLaVA-NeXT-Interleave Li et al. (2024b) LLaVA-NeXT-Interleave-Qwen-7B 7B
@@
Figure 8: Impact of context for plan generation.
Appendix E Hyper-parameters Settings
For closed-source models, results are obtained via API requests during the experimental period from 01/09/2024 to 10/02/2025. The hyper-parameter settings for these API requests include a temperature of 1, top_p of 1, a frequency penalty of 0.2, and a presence penalty of 0.2. All other parameters adhere to the default settings specified by their respective platforms.
Appendix F Automatic Evaluation Metrics
  • BERTScore Zhang et al. (2020): uses contextual embeddings from BERT to evaluate semantic similarity and word overlap between texts.
  • FactVC Liu and Wan (2023): calculates the factual consistency of text with video content by aligning coarse-grained video-text similarity and precision-based fine-grained matching. The values of FactVC range from 0 to 1, and in our experiments, we scale them by 100 to convert them into percentages.
Appendix G Additional Analyses
Impact of Video Context on Summary Generation
We examine the impact of different video context configurations on summary generation, comparing mPLUG-Owl3 with Plan-mPlug-Owl3. Unlike earlier experiments that use the full video as input, here only the first or last 10% or 30% of the video is provided as input. We report results with R2, BERTScore, VideoScore, and FactVC in the full fine-tuning setting.
@@ Appendix G Additional Analyses
Impact of Text Context on Plan Generation
The generation of plan questions in our experiments is influenced by the target sentence and its context. In our main experiments, plan questions are generated based on the target sentence and its preceding summary text (Previous-Context), in line with the original Questions Under Discussion (QUD) requirements Wu et al. (2023a, b). We now assess configurations that generate questions only based on the target sentence (No-Context) or the entire summary (All-Context).
@@
We note that the plan-based method employs an explicit planning component where each sentence is guided by a corresponding question that facilitates fine-grained control over the summary’s style or content. Specifically, after PG produces the plan, we use GPT-o1 to edit it and then feed the edited questions back to SG for the final output. For GPT-o1, which operates in a zero-shot manner, we prepend constraints directly in the prompt. Specifically, GPT-o1 generates an initial summary in one pass and then applies additional prompt-based instructions during a secondary rewriting step to control the output. Both control experiments (Table 7) (Table 8) reveal similar trends: while performance declines for both models, the plan-based method is more robust and controllable.
@@
In the length control experiment (Table 8), R2 scores decline as content is removed, but plan-based model aligns more closely with target compression ratios, producing summaries averaging 100.32 tokens under 60% deletion, while GPT-o1 generates longer summaries (185.28 tokens).
Appendix H Case Study and Error Analysis
For our case study, we randomly select a sample Kübler et al. (2020) from the test split. The analysis in Table 9 reveals differences in summary quality across models, and against the human-written text. Specifically, GPT-o1 often produces concise summaries but at the cost of precision. For example, it incorrectly claims that “data splitting helps control test thresholds,” which is a hallucination — while data splitting ensures a tractable null distribution, it does not explicitly control test thresholds. Furthermore, its summaries frequently oversimplify complex concepts, reducing the depth of explanations and omitting crucial distinctions, such as the role of dependency calibration in the proposed method. Similarly, mPLUG-Owl3 introduces factual inaccuracies, such as stating that data splitting “ensures a reliable null distribution.” This phrasing misleadingly implies that reliability is an inherent property of data splitting, whereas the correct point is that it makes the null distribution tractable rather than necessarily more reliable.
@@ Appendix H Case Study and Error Analysis
Controlled generation experiments also reveal that hallucination issues are further amplified when imposing constraints on readability and length. Under readability control (Table 10), GPT-o1 is more likely to introduce fabricated or misleading content when forced to generate more complex outputs. This occurs because it lacks an explicit mechanism to ensure factual consistency while adapting to varying readability demands. Rather than relying on implicit internal heuristics, Plan-mPlug-Owl3 has an explicit planning mechanism which makes it less likely to introduce unsupported claims. Planning provides an additional layer of control, helping the model maintain factual alignment even as readability demands change.
A similar trend is observed in length control experiments (Table 11). As the compression ratio increases, GPT-o1 struggles to balance conciseness and informativeness, sometimes hallucinating missing details to compensate for omitted content. This suggests that purely instruction-based control (i.e., prompting the model to shorten outputs) does not effectively enforce content retention, leading to greater inconsistencies. In contrast, the plan allows Plan-mPlug-Owl3 to selectively retain essential elements, reducing the risk of generating misleading content; it can also avoid answering deleted questions, to a certain extent.
@@
Reference Modern large-scale kernel-based tests such as maximum mean discrepancy (MMD) and kernelized Stein discrepancy (KSD) optimize kernel hyperparameters on a held-out sample via data splitting to obtain the most powerful test statistics. While data splitting results in a tractable null distribution, it suffers from a reduction in test power due to a smaller test sample size. Inspired by the selective inference framework, we propose an approach that enables learning the hyperparameters and testing on the full sample without data splitting. Our approach can correctly calibrate the test in the presence of such dependency, and yield a test threshold in closed form. At the same significance level, our approach’s test power is empirically larger than that of the data-splitting approach, regardless of its split proportion.
@@
Appendix I LMM-as-Judge Evaluation
To facilitate large-scale comparisons of model outputs, we adopt a method inspired by LLM-as-Judge Liusie et al. (2024); Liu et al. (2024c); Zheng et al. (2024), extending it to use a large multimodal model Chen et al. (2024b). The proposed LMM-based evaluator incorporates both textual and video modalities and assesses the same summary quality dimensions used in our human evaluation. Specifically, we use GPT-o1 as the evaluator, following the hyperparameter settings in Appendix E. To minimize potential bias from prior queries, the conversation history is reset before each evaluation.
@@ Appendix I LMM-as-Judge Evaluation
Compared to fine-tuned models, GPT-o1 assigns the lowest scores to its own responses (see Figure 9). Human-written summaries consistently receive the highest scores and are generally regarded as the best. Aligning with our human evaluations, GPT-o1 also recognizes that the plan-based model outperformed other models. We further conduct paired t-tests to find that human summaries outperform all neural models across all metrics with statistical significance ($p<0.05$). Moreover, the plan-based model demonstrates significantly better performance ($p<0.05$) than other neural models across all metrics except for conciseness. Our results also indicate that although the plan-based method can improve the performance of end-to-end models to some extent, there is a considerable gap between machine-generated and human summaries, which also reflects the challenging nature of our dataset.
Appendix J Prompts Used in Our Study
Quality Control Guidelines Guidelines:
Evaluate each video-text pair to determine whether the text provides a concise and accurate summary of the corresponding video.
Concise: Ensure the text is brief, focused, and free of unnecessary details. Accurate: Verify that the text faithfully represents the video’s content. Make binary judgments (Valid or Invalid) for each pair. If flagged as Invalid, provide a brief justification. Answer:
@@
Prompt for PG model Generate a list of questions for the provided video. Video: {Video} Questions:
@@
Prompt for SG model Generate a summary for the following video based on the plan questions. Video: {Video} Plan Questions: {Questions} Ensure that the generated summary sequentially answers the plan questions. Summary:
@@
Summary Readability Modification Rewrite the following text to further adjust style or detail. Here is the text to be rewritten: {Text} Refine the above text to be more {lay/expert} style. Modified Text:
@@
Summary Length Modification Rewrite the following text to further adjust style or detail. Here is the text to be rewritten: {Text} Shorten the above text by about {10/ 30% / 60%}. Focus on the key points and remove less critical details. Modified Text:
@@
Plan Readability Modification Rewrite the following questions to further adjust style or detail. Here are the questions to be rewritten: 1. {Q1} 2. {Q2} Refine the above questions to be more {lay/expert} style. Modified Questions:
@@
Appendix K Human Evaluation Guidelines
@@
Appendix L Prompt for GPT-o1 to Evaluate Summary Quality

### v4

@@ Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 948878). Lapata acknowledges the support of the UK Engineering and Physical Sciences Research Council (Grant EP/W002876/1).
@@ Appendix D Model Version Details
LLaMA-VID Li et al. (2024c) LLaMA-VID-7B-Full-224-Long-Video 7B
LLaVA-NeXT-Interleave Li et al. (2025) LLaVA-NeXT-Interleave-Qwen-7B 7B
@@
Appendix E Hyper-parameters Settings
For closed-source models, results are obtained via API requests during the experimental period from 01/09/2024 to 10/02/2025. The hyperparameter settings for these API requests include a temperature of 1, top_p of 1, a frequency penalty of 0.2, and a presence penalty of 0.2. All other parameters adhere to the default settings specified by their respective platforms.
Appendix F Automatic Evaluation Metrics
  • BERTScore Zhang et al. (2020): uses contextual embeddings from BERT to evaluate semantic similarity between texts.
  • FactVC Liu and Wan (2023): calculates the factual consistency of text with video content by aligning coarse-grained video-text similarity and precision-based fine-grained matching. The original values of FactVC range from 0 to 1, and in our experiments, we scale them by 100 to convert them into percentages.
Appendix G Plans Quality Validation
To validate the quality of the silver-standard plans generated by GPT-o1, we conduct a manual evaluation on 100 randomly selected samples. The evaluation is carried out by the same annotators involved in our human evaluation setup. Each annotator is asked to make a binary judgment on whether the generated plan question satisfied two validity criteria: (1) Local Coherence: The question is well-formed and semantically related to the summary; and (2) QUD-Alignment: Each sentence in the summary could plausibly serve as an answer to the question, consistent with the QUD framework.
We observe strong inter-annotator agreement (Fleiss’ $\kappa=0.853$), indicating a high degree of consistency in decisions. In addition to this, we perform a manual error analysis to screen for systematic biases or recurrent flaws, such as overly generic phrasing, hallucinated entities, or structural redundancy. No such patterns are observed.
Appendix H Impact of Video Context on Summary Generation
We examine the impact of different video context configurations on summary generation, comparing mPLUG-Owl3 with Plan-mPlug-Owl3. Unlike earlier experiments that use the full video as input, here only the first or last 10% or 30% of the video is provided as input. We report results in the full fine-tuning setting.
@@ Appendix H Impact of Video Context on Summary Generation
Appendix I Impact of Text Context on Plan Generation
The generation of plan questions in our experiments is influenced by the target sentence and its context. In our main experiments, plan questions are generated based on the target sentence and its preceding summary text (Previous-Context), in line with the original Question Under Discussion (QUD) requirements Wu et al. (2023a, b); Liu et al. (2025). We now assess configurations that generate questions only based on the target sentence (No-Context) or the entire summary (All-Context).
Figure 8: Impact of text context for plan generation.
@@ Appendix I Impact of Text Context on Plan Generation
We note that the plan-based method employs an explicit planning component where each sentence is guided by a corresponding question that facilitates fine-grained control over the summary’s style or content. Specifically, after PG produces the plan, we use GPT-o1 to edit it and then feed the edited questions back to SG for the final output. For GPT-o1, which operates in a zero-shot manner, we prepend constraints directly in the prompt. Specifically, GPT-o1 generates an initial summary in one pass and then applies additional prompt-based instructions during a secondary rewriting step to control the output. Both control experiments (Table 9 and Table 10) reveal similar trends: While performance declines for both models, the plan-based method is more robust and controllable.
@@
In the length control experiment (Table 10), R2 scores decline as content is removed, but the plan-based model aligns more closely with target compression ratios, producing summaries averaging 100.32 tokens under 60% deletion, while GPT-o1 generates longer summaries (185.28 tokens).
Appendix K Case Study and Error Analysis
For our case study, we randomly select a sample Kübler et al. (2020) from the test split. The analysis in Table 11 reveals differences in summary quality across models and against the human-written text. Specifically, GPT-o1 often produces concise summaries but at the cost of precision. For example, it incorrectly claims that “data splitting helps control test thresholds,” which is a hallucination — while data splitting ensures a tractable null distribution, it does not explicitly control test thresholds. Furthermore, its summaries frequently oversimplify complex concepts, reducing the depth of explanations and omitting crucial distinctions, such as the role of dependency calibration in the proposed method. Similarly, mPLUG-Owl3 introduces factual inaccuracies, such as stating that data splitting “ensures a reliable null distribution.” This phrasing misleadingly implies that reliability is an inherent property of data splitting, whereas the correct point is that it makes the null distribution tractable rather than necessarily more reliable.
@@
Controlled generation experiments reveal that hallucination issues are further amplified when imposing constraints on readability and length. Under readability control (Table 12), GPT-o1 is more likely to introduce fabricated or misleading content when forced to generate more complex outputs. This occurs because it lacks an explicit mechanism to ensure factual consistency while adapting to varying readability demands. Rather than relying on implicit internal heuristics, Plan-mPlug-Owl3 has an explicit planning mechanism which makes it less likely to introduce unsupported claims. Planning provides an additional layer of control, helping the model maintain factual alignment even as readability demands change. A similar trend is observed in length control experiments (Table 13). As the compression ratio increases, GPT-o1 struggles to balance conciseness and informativeness, sometimes hallucinating missing details to compensate for omitted content. This suggests that purely instruction-based control (i.e., prompting the model to shorten outputs) does not effectively enforce content retention, leading to greater inconsistencies. In contrast, the plan allows Plan-mPlug-Owl3 to selectively retain essential elements, reducing the risk of generating misleading content; it can also avoid answering deleted questions, to a certain extent.
@@
Reference Modern large-scale kernel-based tests such as maximum mean discrepancy (MMD) and kernelized Stein discrepancy (KSD) optimize kernel hyperparameters on a held-out sample via data splitting to obtain the most powerful test statistics. While data splitting results in a tractable null distribution, it suffers from a reduction in test power due to a smaller test sample size. Inspired by the selective inference framework, we propose an approach that enables learning the hyperparameters and testing on the full sample without data splitting. Our approach can correctly calibrate the test in the presence of such dependency and yield a test threshold in closed form. At the same significance level, our approach’s test power is empirically larger than that of the data-splitting approach, regardless of its split proportion.
@@
Appendix L LMM-as-Judge Evaluation
To facilitate large-scale comparisons of model outputs, we adopt a method inspired by LLM-as-Judge Liusie et al. (2024); Liu et al. (2024c); Zheng et al. (2024); Liu et al. (2025), extending it to use a large multimodal model Chen et al. (2024b). The proposed LMM-based evaluator incorporates both textual and video modalities and assesses the same summary quality dimensions used in our human evaluation for trend-level comparisons. Specifically, we use GPT-o1 as the evaluator, following the hyperparameter settings in Appendix E. To minimize potential bias from prior queries, the conversation history is reset before each evaluation. The instructions for the LMM-as-Judge evaluation are provided in Appendix O.
@@
Compared to fine-tuned models, GPT-o1 assigns the lowest scores to its own responses (see Figure 9). Human-written summaries consistently receive the highest scores and are generally regarded as the best. Aligning with our human evaluations, GPT-o1 also recognizes that the plan-based model outperformed other models. We further conduct paired t-tests to find that human summaries outperform all neural models across all metrics with statistical significance ($p<0.05$). Moreover, the plan-based model demonstrates significantly better performance ($p<0.05$) than other neural models across all metrics except for conciseness. Our results also indicate that although the plan-based method can improve the performance of end-to-end models to some extent, there is still a gap between machine-generated and human summaries, which also reflects the challenging nature of our dataset.
Appendix M Prompts Used in Our Study
Quality Control Guidelines Evaluate each video-text pair to determine whether the text provides a concise and accurate summary of the corresponding video. Concise: Ensure the text is brief, focused, and free of unnecessary details. Accurate: Verify that the text faithfully represents the video’s content. Make binary judgments (Valid or Invalid) for each pair. If flagged as Invalid, provide a brief justification. Answer:
@@
Prompt for PG model Generate a list of questions for the provided {Video/Audio/Transcript...}. Content: {Video/Audio/Transcript...} Questions:
@@
Prompt for SG model Generate a summary for the following {Video/Audio/Transcript...} based on the plan questions. Content: {Video/Audio/Transcript...}. Plan Questions: {Questions} Ensure that the generated summary sequentially answers the plan questions. Summary:
@@
Summary Readability Modification Rewrite the following text to further adjust the style or detail. Here is the text to be rewritten: {Text} Refine the above text to be more {lay/expert} style. Modified Text:
@@
Summary Length Modification Rewrite the following text to further adjust the style or detail. Here is the text to be rewritten: {Text} Shorten the above text by about {10/ 30% / 60%}. Focus on the key points and remove less critical details. Modified Text:
@@
Plan Readability Modification Rewrite the following questions to further adjust the style or detail. Here are the questions to be rewritten: 1. {Q1} 2. {Q2} Refine the above questions to be more {lay/expert} style. Modified Questions:
@@
Appendix N Human Evaluation Guidelines
@@
Appendix O Prompt for GPT-o1 to Evaluate Summary Quality